Problem 2: Fill in the following table: u (kJ/kg) State p(atm) OPEL The temperature at the beginning of the compression proce fill in the table please
Answer #1

Que Air-standart Orro Cycle - Networkourpur Carca) vac osoner 3 pvc V=C Given V V = 0.001 me Iz 140°c = 273 +140 = 443K Pg Vyx-1 C = 1.4 ( År sraudand cycle) 0.4 1-1-6) Z=14 ( Aur creandard % = 1- (6 jou = 0151164 | Piu Pivit % = 51.1864% we know for0.546x106 we know from the first law of thermodynamice EQ = du tow for process 1-2 r 1-2 a dichatic proveis (compressing 88,0& Wher = 8Wzy t&Wiz & Wher = 476.974 - 80 813.669 swner = 1611301 Kylkg (ND) ner = 161.305 Kylkg. (WD) ner = Prean x V efrect* power per cycle & power = (Wilhet x mpreg. Cycle - 161.305 X 8.5UB X106 Power 1.3786 x 103 Kylagde Pomep = 201.63 1 mpa Vac